OpenStudy (anonymous):

I have five questions i need answered. Its solving systems! Can some one help?

OpenStudy (texaschic101):

whats the problem ?

OpenStudy (anonymous):

i have to do number 82-86 and i cant figure any of them out i did 81 but they were really hard

OpenStudy (anonymous):

please please please help

OpenStudy (anonymous):

pleeeeeeease help im begging

OpenStudy (texaschic101):

I am not too good at these.... y = x^2 + 2x - 8 ....(y = x + 4) now we sub x + 4 = x^2 + 2x - 8 x^2 + 2x - x - 8 - 4 = 0 x^2 + x - 12 = 0 (x - 3)(x + 4) = 0 x - 3 = 0 x = 3 x + 4 = 0 x = -4 x = 3 and x = -4 now I am guessing you just sub both of those answers into y = x + 4, and solve for y ??

OpenStudy (anonymous):

see if it works

OpenStudy (anonymous):

i got that one right! :D yay!

OpenStudy (texaschic101):

when x = 3, y = x + 4 y = 3 + 4 y = 7 and when x = -4, y = x + 4 y = -4 + 4 y = 0 check.. y = x^2 + 2x - 8 7 = 3^2 + 2(3) - 8 7 = 9 + 6 - 8 7 = 15 - 8 7 = 7 (correct) y = x^2 + 2x - 8 0 = -4^2 + 2(-4) - 8 0 = 16 - 8 - 8 0 = 0 (correct)

OpenStudy (anonymous):

can you help me especiially with number 84-86 cause i dont understand those at alllll

OpenStudy (texaschic101):

y = x^2 - 3x - 7 y = -x^2 - x + 5 set them equal to each other x^2 - 3x - 7 = -x^2 - x + 5 x^2 + x^2 - 3x + x - 7 - 5 = 0 2x^2 - 2x - 12 = 0 2(x + 2)(x - 3) x + 2 = 0 x = -2 x - 3 = 0 x = 3 now pick an equation and sub in your x answers for y y = x^2 - 3x - 7 y = -2^2 - 3(-2) - 7 y = 4 + 6 - 7 y = 10 - 7 y = 3 y = x^2 - 3x - 7 y = 3^2 - 3(3) - 7 y = 9 - 9 - 7 y = -7

OpenStudy (anonymous):

can you do 86? im so confused!!! ive never seen fraction in them. my teacher didnt go over it

OpenStudy (texaschic101):

y = x^2 - 2x - 1 y = 3/4x^2 + x - 6 set them equal x^2 - 2x - 1 = 3/4x^2 + x - 6 4/4x^2 - 3/4x^2 - 2x - x - 1 + 6 = 0 1/4x^2 - 3x + 5 = 0 Oh crap....I am sorry. I do not know what to do next :( I am probably totally wrong on this...sorry again

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

85?

OpenStudy (anonymous):

i figured them out...for future reference..you divide 1/4 by all 3 numbers to get answers!